# Resummed Distribution Functions: Making Perturbation Theory Positive and Normalized

### v1

@@ 7 Discussion and Conclusions
There are a variety of settings where the RDF can be applied. In cases where it is either difficult or impossible to resum a fixed-order calculation as is the case with simultaneous observables Larkoski:2014tva ; Lustermans:2019gxu , such as Energy Flow Polynomials Komiske:2017aww ; Cal:2022fnm , the RDF is a method for at constraining and reasonably guessing the all-orders structure of the cross section. The RDF requires no additional “physics” knowledge to be applied (unlike a genuine resummation) other than that the theory predicts valid probabilities. The built-in method of varying higher-order terms can be used to easily understand theoretical uncertainties, either qualitatively through random variations as was performed in Secs. 4 and 5, or quantitatively with nuisance parameters as was performed in Sec. 6. In Sec. 6, we showed the utility of the RDF for phenomenology studies by extracting the strong coupling constant from an RDF-matched calculation to thrust. We might further imagine fitting the RDF to a diverse set of other observables in which resummation effects or theoretical uncertainties can be important, such as $p_{T}$ tails, electroweak fits, or other event shapes, such as C-parameter Hoang:2015hka .
@@ A.1 Supplementary Plots for Thrust Fits
Figure 11: The same as the right panel of Fig. 10, but with a Gaussian prior scale on $g_{mn}$ of $\sigma=0.5$ (left) and $\sigma=2.0$ (right) instead of $\sigma=1.0$
In Fig. 11, we show the analogue of Fig. 10, but with $\sigma=0.5$ and $\sigma=2.0$ as shown in Table 1. While the $\mathcal{O}(\alpha_{s}^{2})$ and $\mathcal{O}(\alpha_{s}^{3})$ are roughly consistent between all the prior choices, the $\mathcal{O}(\alpha_{s}^{1})$ calculation fit is not. They are much greater than the expected value of $\alpha_{s}\approx 0.118$, and in fact, the $\sigma=2.0$ fit reaches the edge of our fit range.
Figure 12: The same as the right panel of Fig. 10, but with a completely flat prior
@@ A.2 Training Plots
In this appendix, we show an example of the numeric matching procedure. The purpose of this appendix is primarily to give a semi-pedagogical taste of the numeric matching for readers unfamiliar with this machine-learning style minimization, by attaching some real numbers to each of the parameter objects. We elect only to show the $\mathcal{O}(\alpha_{s}^{1})$ exponential matching, corresponding to the top row of Fig. 3. This is for several reasons: first and foremost, the actual values of the $g$ matrices are not necessarily informative, except in the toys where we can check if they match the analytic expectations. One way to see this is that the $T$-parameters can be slightly degenerate with the $g$ parameters, making both more complex to interpret. Second, the $g$ matrices can be rather large beyond $\mathcal{O}(\alpha_{s}^{1})$, and there is little information to gain by staring at a large number of minimization curves.
Refer to captionRefer to caption
@@ B.1 Reroll initialization procedure
which decreases linearly to zero over the entire initialization procedure, and further decreases each time a new best loss os found (the number of times is given by “counts”). We run 1024 of these initialization threads in parallel, and we take the set of parameters with the lowest loss at the end as our initialization for gradient descent.

### v2

@@ 7 Discussion and Conclusions
There are a variety of settings where the RDF can be applied. In cases where it is either difficult or impossible to resum a fixed-order calculation as is the case with simultaneous observables Larkoski:2014tva ; Lustermans:2019gxu , such as Energy Flow Polynomials Komiske:2017aww ; Cal:2022fnm , the RDF is a method for constraining and reasonably guessing the all-orders structure of the cross section. The RDF requires no additional “physics” knowledge to be applied (unlike a genuine resummation) other than that the theory predicts valid probabilities. The built-in method of varying higher-order terms can be used to easily understand theoretical uncertainties, either qualitatively through random variations as was performed in Secs. 4 and 5, or quantitatively with nuisance parameters as was performed in Sec. 6. In Sec. 6, we showed the utility of the RDF for phenomenology studies by extracting the strong coupling constant from an RDF-matched calculation to thrust. We might further imagine fitting the RDF to a diverse set of other observables in which resummation effects or theoretical uncertainties can be important, such as $p_{T}$ tails, electroweak fits, or other event shapes, such as C-parameter Hoang:2015hka . It would also be interesting to explore bases for $g$ beyond polynomial (or, to potentially replace $g$ with neural networks), which might be more suited to observables with less trivial fixed-order expansions like C-parameter.
@@ A.1 Supplementary Plots for Thrust Fits
Figure 11: The same as the right panel of Fig. 10, but with a Gaussian prior scale on $g_{mn}$ of $\sigma=0.5$ (left) and $\sigma=2.0$ (right) instead of $\sigma=1.0$.
In Fig. 11, we show the analogue of Fig. 10, but with $\sigma=0.5$ and $\sigma=2.0$ as shown in Table 1. While the $\mathcal{O}(\alpha_{s}^{2})$ and $\mathcal{O}(\alpha_{s}^{3})$ results are roughly consistent between all the prior choices, the $\mathcal{O}(\alpha_{s}^{1})$ calculation fit is not. They are much greater than the expected value of $\alpha_{s}\approx 0.118$, and in fact, the $\sigma=2.0$ fit reaches the edge of our fit range.
Figure 12: The same as the right panel of Fig. 10, but with a completely flat prior.
@@ A.2 Training Plots
In this appendix, we show an example of the numeric matching procedure. The purpose of this appendix is primarily to give a semi-pedagogical taste of the numeric matching for readers unfamiliar with this machine-learning style minimization by attaching some real numbers to each of the parameter objects. We elect only to show the $\mathcal{O}(\alpha_{s}^{1})$ exponential matching, corresponding to the top row of Fig. 3. This is for several reasons: first and foremost, the actual values of the $g$ matrices are not necessarily informative, except in the toys where we can check if they match the analytic expectations. One way to see this is that the $T$-parameters can be slightly degenerate with the $g$ parameters, making both more complex to interpret. Second, the $g$ matrices can be rather large beyond $\mathcal{O}(\alpha_{s}^{1})$, and there is little information to gain by staring at a large number of minimization curves.
@@ B.1 Reroll initialization procedure
which decreases linearly to zero over the entire initialization procedure, and further decreases each time a new best loss is found (the number of times is given by “counts”). We run 1024 of these initialization threads in parallel, and we take the set of parameters with the lowest loss at the end as our initialization for gradient descent.